213) SR2 + SR11 =
224) SR(6)*(SR6 – SR2) =
265) (SR5 + 3)(SR 2 – 8) =
: Simplify or solve if it is an equation. SR is square root

Answer by stanbon(75887)

213) SR2 + SR11 =
No way to simplify
----------------------------------------
224) SR(6)*(SR6 – SR2) = 6 - SR(12) = 6 - 2SR(3)
---------------------------------------
265) (SR5 + 3)(SR 2 – 8)
= SR(10) -8SR(5) + 3SR(2) - 24
--------------------------------------------
Cheers,
Stan H.